# Non-variational Quantum Combinatorial Optimisation

### v1

@@ III-A Binary Mixer
III-B Integer Mixer
Consider a combinatorial problem consisting of $n$ integer decision variables ${x_{j}\in\{0,1,...,k-1\}}$, where solutions are characterised by vectors ${\bm{x}=(x_{1},x_{2},...,x_{n})}$ and the space of feasible solutions $S$ contains all such vectors, such that $N=k^{n}$. Each integer variable can be encoded in the computational basis state of a sub-register of $\lceil\log_{2}k\rceil$ qubits. The equal superposition state is efficiently prepared over an $n\lceil\log_{2}k\rceil$ qubit register containing $n$ sub-registers,
|$\ket{s}=\left(U_{k}\ket{0}^{\otimes\lceil\log_{2}k\rceil}\right)^{\otimes n}=\ket{k}^{\otimes n}=\frac{1}{\sqrt{k^{n}}}\sum_{\bm{x}\in S}\ket{\bm{x}},$ (13)
where |$\ket{k}$ is the equal superposition over the first $k$ computational basis states of a sub-register. A circuit implementation of the unitary $U_{k}$ which prepares |$\ket{k}$ is described in [21].
@@ III-B Integer Mixer
where $K_{k}$ is the adjacency matrix of a complete graph connecting the first $k$ computational basis states within a sub-register, ${K_{k}=k\outerproduct{k}{k}-\bbold{1}}$.
@@
III-C Permutation Mixer
Consider a combinatorial problem consisting of $n$ integer decision variables ${x_{j}\in\{0,1,...,n-1\}}$, where solutions are characterised by vectors ${\bm{x}=(x_{1},x_{2},...,x_{n})}$ and the space of feasible solutions $S$ contains all such vectors without repeating elements, such that $N=n!$. As with the integer problems, each integer variable can be encoded in the computational basis state of a sub-register of $\lceil\log_{2}n\rceil$ qubits. A method which can be used to generate the equal superposition state is to first produce an equal superposition over the first $N$ computational basis states of a separate $\lceil\log_{2}N\rceil$ qubit register, followed by the use of an appropriate indexing algorithm as discussed in more detail in [21]. Note that the method presented in [22] produces permutations stored with a non-binary (one-hot) encoding, rather than a binary encoding, so it is not directly applicable (and requires many more qubits).
@@ III-C Permutation Mixer
where $\text{SWAP}_{i,j}$ is defined as the permutation matrix associated with swapping the states of the $i^{\text{th}}$ and $j^{\text{th}}$ sub-registers (applying identity to the remaining sub-registers).
@@ IV Simulation Results
This section introduces several problems and provides simulation results for the non-variational QWOA applied to each of them. For this purpose, a single random problem instance has been generated for each, with the details provided in [21]. In addition, each of the generated problems are shown to satisfy the necessary condition in (8), where the result is derived analytically for weighted maxcut, and is demonstrated via statistical sampling for the other problems. Note that the approximation ratio, defined as,
@@ IV-A Weighted Maxcut
Figure 3: An analysis of objective function values for the $n=18$ weighted maxcut problem, as distributed over the binary/hypercube mixing graph.
With the binary mixer, $\beta=0.4$ and $p=10$, the optimal amplified state |$\ket{\gamma,t}$ for an $n=18$ problem instance is determined by gradient ascent with initial parameter values $\gamma=1$ and $t=0.1$. The optimal values are $\gamma=1.6031$ and $t=0.4008$, for which the amplified state is analysed in Fig. 4 and Fig. 5, demonstrating that the non-variational QWOA solves this particular instance.
Figure 4: Measurement probability of the optimal solution throughout the $10$ iterations preparing the amplified state |$\ket{\gamma,t}$.
Figure 5: Distributions for approximation ratio as measured from the initial equal superposition state, |$\ket{s}$, and the amplified state |$\ket{\gamma,t}$.
@@ IV-B Maximum Independent Set
Figure 6: An analysis of objective function values with $\bm{\lambda}=\{1.5,0\}$ for the $n=18$ maximum independent set problem, as distributed over the binary/hypercube mixing graph. The points and shading show the mean and $\pm$ a single standard deviation from 200 equally spaced bins (where the bins were not empty).
With the binary mixer, $\beta=\frac{1}{p}$ and $p=10$, the optimal amplified state |$\ket{\gamma,t}_{\bm{\lambda}_{T}}$ for an $n=18$ problem instance is determined by gradient ascent with initial parameter values $\gamma=1$, $t=0.1$ and ${\bm{\lambda}_{T}=\bm{\lambda}_{F}=\{1.5,0\}}$. The optimal values are $\gamma=3.3647$, $t=0.6156$, and ${\bm{\lambda}_{T}=\{1.1094,0.4747\}}$, for which the amplified state is analysed in Fig. 7 and Fig. 8, demonstrating that the non-variational QWOA solves this particular instance, preferentially amplifying the measurement probability of the two maximum independent sets.
Figure 7: Measurement probability of the two optimal solutions throughout the $10$ iterations preparing the amplified state |$\ket{\gamma,t}_{\bm{\lambda}_{T}}$.
Figure 8: Distributions for $f(\bm{x})_{\bm{\lambda}_{F}}$ as measured from the initial equal superposition state, |$\ket{s}$, and the amplified state |$\ket{\gamma,t}_{\bm{\lambda}_{T}}$.
@@
In [21] an additional adjustment is made to the objective function to improve its adherence to the condition in (8),
@@
Figure 9: An analysis of objective function values for the $n=12$, $k=3$ k-means clustering problem, as distributed over the integer mixing graph. The points and shading show the mean and $\pm$ a single standard deviation from 200 equally spaced bins (where the bins were not empty).
With the integer mixer, $\beta=0.4$ and $p=10$, the optimal amplified state |$\ket{\gamma,t}$ for an $n=12$, $k=3$ problem instance is determined by gradient descent with initial parameter values $\gamma=1$ and $t=0.1$. The optimal values are $\gamma=1.4174$ and $t=0.2434$, for which the amplified state is analysed in Fig. 10 and Fig. 11, demonstrating that the non-variational QWOA solves this particular instance.
Figure 10: Measurement probability of the optimal solution throughout the $10$ iterations preparing the amplified state |$\ket{\gamma,t}$.
Figure 11: Distributions for approximation ratio as measured from the initial equal superposition state, |$\ket{s}$, and the amplified state |$\ket{\gamma,t}$.
@@ IV-D The Capacitated Facility Location Problem
Figure 12: An analysis of objective function values $f(\bm{x})_{\bm{\lambda}_{T}}$ for the $n=12$, $k=3$ capacitated facility location problem, as distributed over the integer mixing graph. The points and shading show the mean and $\pm$ a single standard deviation from 200 equally spaced bins (where the bins were not empty).
With the integer mixer, $\beta=\frac{1}{p}$ and $p=20$, the optimal amplified state |$\ket{\gamma,t}_{\bm{\lambda}_{T}}$ for an $n=12$, $k=3$ problem instance is determined by gradient ascent with initial parameter values $\gamma=1$, $t=0.1$ and ${\bm{\lambda}_{T}=\bm{\lambda}_{F}=\{1,1,0\}}$. The optimal values are $\gamma=2.6246$, $t=0.2798$ and $\bm{\lambda}_{T}=\{0.8880,0.4985,0.1731\}$, for which the amplified state is analysed in Fig. 13 and Fig. 14, demonstrating that the non-variational QWOA solves this particular instance.
Figure 13: Measurement probability of the optimal solution throughout the $20$ iterations preparing the amplified state |$\ket{\gamma,t}_{\bm{\lambda}_{T}}$.
Figure 14: Distributions for $f(\bm{x})_{\bm{\lambda}_{F}}$ as measured from the initial equal superposition state, |$\ket{s}$, and the amplified state |$\ket{\gamma,t}_{\bm{\lambda}_{T}}$.
@@ IV-E The Quadratic Assignment Problem
Figure 15: An analysis of objective function values $f(\bm{x})$ for the $n=9$ QAP, as distributed over the permutation mixing graph. The points and shading show the mean and $\pm$ a single standard deviation from 200 equally spaced bins (where the bins were not empty).
With the permutation mixer, $\beta=0.4$ and $p=20$, the optimal amplified state |$\ket{\gamma,t}$ for an $n=9$ problem instance is determined by gradient descent with initial parameter values $\gamma=1$ and $t=0.1$. The optimal values are $\gamma=1.2818$ and $t=0.1219$, for which the amplified state is analysed in Fig. 16 and Fig. 17, demonstrating that the non-variational QWOA solves this particular instance.
Figure 16: Measurement probability of the optimal solution throughout the $20$ iterations preparing the amplified state |$\ket{\gamma,t}$.
Figure 17: Distributions for $f(\bm{x})$ as measured from the initial equal superposition state, |$\ket{s}$, and the amplified state |$\ket{\gamma,t}$.
@@ V Conclusion
We characterise the algorithm as non-variational because, given a number of iterations $p$, the amplified state is determined by a small set of parameters $\{\sigma,\beta,\gamma,t\}$. The number of iterations necessary for a particular problem and problem size can be determined via benchmarking, or may be limited by achievable circuit depths for a particular quantum computing device. The approximate standard deviation $\sigma$ of $f(\bm{x})$ can be easily determined via small scale random sampling. The user-specified parameter $\beta$ only needs to be determined once for a particular problem type. Finally, the two parameters $\gamma$ and $t$ are relatively consistent across problem instances, but they can be fine-tuned via 2-dimensional local optimization to further improve the performance of the amplified state |$\ket{\gamma,t}$.
In order to quantify the algorithm’s average-case and worst-case performance, an important future area of research will involve detailed and large scale benchmarking for various problems. We also plan to explore the application of the non-variational QWOA to the capacitated vehicle routing problem and the multiple traveling salesman problem.

### v2

@@ III-A Binary Mixer (Hypercube Graph)
III-B Integer Mixer (Hamming Graph)
Consider a combinatorial problem consisting of $n$ integer decision variables ${x_{j}\in\{0,1,...,k-1\}}$, where solutions are characterised by vectors ${\bm{x}=(x_{1},x_{2},...,x_{n})}$ and the space of feasible solutions $S$ contains all such vectors, such that ${N=k^{n}}$. Each integer variable can be encoded in the computational basis state of a sub-register of $m=\lceil\log_{2}k\rceil$ qubits for a binary encoding or $m=k$ qubits for a one-hot encoding. The equal superposition state is efficiently prepared over an $nm$ qubit register containing $n$ sub-registers,
|$\ket{s}=\left(U_{k}\ket{0}^{\otimes m}\right)^{\otimes n}=\ket{k}^{\otimes n}=\frac{1}{\sqrt{k^{n}}}\sum_{\bm{x}\in S}\ket{\bm{x}},$ (13)
where |$\ket{k}$ is the equal superposition over the $k$ utilized computational basis states of a sub-register (those which directly encode each of the possible decision variable values). Each circuit implementation (one-hot and binary) of the unitary $U_{k}$ which prepares |$\ket{k}$ is described in [21].
@@ III-B Integer Mixer (Hamming Graph)
where $K_{k}$ is the adjacency matrix of a complete graph connecting the $k$ utilized computational basis states within a sub-register, ${K_{k}=k\outerproduct{k}{k}-\bbold{1}}$. This mixing graph (the subgraph connecting feasible solutions) is also known as a Hamming graph $H(n,k)$, and is the Cartesian product of $n$ complete graphs $K_{k}$.
@@
III-C Permutation Mixer (Transposition Graph)
Consider a combinatorial problem consisting of $n$ integer decision variables ${x_{j}\in\{0,1,...,n-1\}}$, where solutions are characterised by vectors ${\bm{x}=(x_{1},x_{2},...,x_{n})}$ and the space of feasible solutions $S$ contains all such vectors without repeating elements, such that $N=n!$. As with the integer problems, each integer variable can be encoded in the computational basis state of a sub-register of $m=\lceil\log_{2}n\rceil$ or $m=k$ qubits. A recursive method for generating the equal superposition over feasible solution states is presented in [21], which has gate complexity of order $n^{3}$, and works for both the binary and one-hot encodings. An alternative method is presented in [23], effective for just the one-hot encoding.
@@ III-C Permutation Mixer (Transposition Graph)
where $\text{SWAP}_{i,j}$ is defined as the permutation matrix associated with swapping the states of the $i^{\text{th}}$ and $j^{\text{th}}$ sub-registers (applying identity to the remaining sub-registers). This mixing graph (the subgraph connecting feasible solutions) is a transposition graph.
@@ IV Simulation Results
This section introduces several problems and provides simulation results for the non-variational QWOA applied to each of them. For this purpose, a single random problem instance has been generated for each, with the details provided in [21]. In addition, each of the generated problems are shown to satisfy the necessary condition in (8), where the result is derived analytically for weighted maxcut, and is demonstrated via statistical sampling for the other problems.
During physical implementation, derivative-free optimisation methods such as Bayesian optimisation or CMA-ES, are likely to be more effective than gradient-based methods, as they are robust to uncertainties in the approximated values of the function to be optimised. This allows them to be effective with limited shots (repeated state-preparation and measurements). However, for the following results, parameter values have all been determined via the BFGS algorithm, to optimise for expectation value of $f(\bm{x})$, initialised with $\gamma=1$, $t=0.1$ and $\beta=\frac{1}{p}$ (and $\bm{\lambda}_{T}=\bm{\lambda}_{F}$ where relevant).
Note that the approximation ratio, defined as,
@@ IV-A Weighted Maxcut
With the binary mixer and $p=10$, the optimal amplified state |$\ket{\gamma,t,\beta}$ for an $n=18$ problem instance is given by $\gamma=2.4340$, $t=0.4517$ and $\beta=0.2844$ and is analysed in Fig. 4 and Fig. 5, demonstrating that the non-variational QWOA solves this particular instance.
Figure 3: Maxcut: An analysis of objective function values, as distributed over the binary/hypercube mixing graph.
Figure 4: Maxcut: Measurement probability of the optimal solution throughout the $10$ iterations preparing the amplified state |$\ket{\gamma,t}$.
Figure 5: Maxcut: Distributions for approximation ratio as measured from the initial equal superposition state, |$\ket{s}$, and the amplified state |$\ket{\gamma,t}$.
@@ IV-B Maximum Independent Set
Figure 6: Maximum independent set: An analysis of objective function values with $\bm{\lambda}_{F}=\{1.5,0\}$, as distributed over the binary/hypercube mixing graph. The points and shading show the mean and $\pm$ a single standard deviation from 200 equally spaced bins (where the bins were not empty).
With the binary mixer, $p=10$ and $\bm{\lambda}_{F}=\{1.5,0\}$, the optimal amplified state |$\ket{\gamma,t,\beta}_{\bm{\lambda}_{T}}$ for an $n=18$ problem instance is given by parameter values, $\gamma=3.0098$, $t=0.5724$, $\beta=0.1722$ and ${\bm{\lambda}_{T}=(1.0370,0.5235)}$ and is analysed in Fig. 7 and Fig. 8, demonstrating that the non-variational QWOA solves this particular instance, preferentially amplifying the measurement probability of the two maximum independent sets.
Figure 7: Maximum independent set: Measurement probability of the two optimal solutions throughout the $10$ iterations preparing the amplified state |$\ket{\gamma,t,\beta}_{\bm{\lambda}_{T}}$.
Figure 8: Maximum independent set: Distributions for $f(\bm{x})_{\bm{\lambda}_{F}}$ as measured from the initial equal superposition state, |$\ket{s}$, and the amplified state |$\ket{\gamma,t,\beta}_{\bm{\lambda}_{T}}$.
@@
An additional adjustment is made to the objective function to improve its adherence to the condition in (8),
@@
Figure 9: k-means clustering: An analysis of objective function values, as distributed over the integer mixing graph. The points and shading show the mean and $\pm$ a single standard deviation from 200 equally spaced bins (where the bins were not empty).
With the integer mixer and $p=10$, the optimal amplified state |$\ket{\gamma,t,\beta}$ for an $n=12$, $k=3$ problem instance is given by parameter values $\gamma=1.5345$, $t=0.2483$ and $\beta=0.3441$ and is analysed in Fig. 10, demonstrating that the non-variational QWOA solves this particular instance.
Figure 10: k-means clustering: Measurement probability of the optimal solution throughout the $10$ iterations preparing the amplified state |$\ket{\gamma,t}$.
@@ IV-D The Capacitated Facility Location Problem
Figure 11: Capacitated facility location problem: An analysis of objective function values $f(\bm{x})_{\bm{\lambda}_{T}}$ as distributed over the integer mixing graph. The points and shading show the mean and $\pm$ a single standard deviation from 200 equally spaced bins (where the bins were not empty).
With the integer mixer, $p=20$ and ${\bm{\lambda}_{F}=(1,1,0)}$, the optimal amplified state |$\ket{\gamma,t,\beta}_{\bm{\lambda}_{T}}$ for an $n=12$, $k=3$ problem instance is given by $\gamma=2.5732$, $t=0.2756$, $\beta=0.0593$ and ${\bm{\lambda}_{T}=(0.8966,0.4996,0.1732)}$ and is analysed in Fig. 12 and Fig. 13, demonstrating that the non-variational QWOA solves this particular instance.
Figure 12: Capacitated facility location problem: Measurement probability of the optimal solution throughout the $20$ iterations preparing the amplified state |$\ket{\gamma,t}_{\bm{\lambda}_{T}}$.
Figure 13: Capacitated facility location problem: Distributions for $f(\bm{x})_{\bm{\lambda}_{F}}$ as measured from the initial equal superposition state, |$\ket{s}$, and the amplified state |$\ket{\gamma,t,\beta}_{\bm{\lambda}_{T}}$.
@@ IV-E The Quadratic Assignment Problem
Figure 14: The quadratic assignment problem: An analysis of objective function values $f(\bm{x})$ as distributed over the permutation mixing graph. The points and shading show the mean and $\pm$ a single standard deviation from 200 equally spaced bins (where the bins were not empty).
With the permutation mixer and $p=20$, the optimal amplified state |$\ket{\gamma,t,\beta}$ for an $n=9$ problem instance is given by $\gamma=1.2636$, $t=0.1219$ and $\beta=0.4167$ and is analysed in Fig. 15, demonstrating that the non-variational QWOA solves this particular instance.
Figure 15: The quadratic assignment problem: Measurement probability of the optimal solution throughout the $20$ iterations preparing the amplified state |$\ket{\gamma,t}$.
@@ V Conclusion
We characterise the algorithm as non-variational because, given a number of iterations $p$, the amplified state is determined by a small set of parameters $\{\sigma,\beta,\gamma,t\}$. The number of iterations necessary for a particular problem and problem size can be determined via benchmarking, or may be limited by achievable circuit depths for a particular quantum computing device. The approximate standard deviation $\sigma$ of $f(\bm{x})$ can be easily determined via small scale random sampling. The user-specified parameters $\gamma$, $t$ and $\beta$, produce significant amplification of optimal and near-optimal solutions across a wide range of values. In addition, the optimal values are relatively consistent across problem instances, and can also be found (or improved) via 3-dimensional local optimisation.
The number of state-preparation and measurements required to perform this local optimisation is fixed, independent of iteration count or problem size. Hence, one important component of future work will involve the deployment of gradient-free optimisation methods, such as Bayesian optimisation and CMA-ES, in order to demonstrate the ability to find/improve parameter values within a fixed and small number of total state preparations/measurements ($<10000$).
In order to quantify the algorithm’s average-case performance, and demonstrate advantage relative to classical heuristics, such as local search based methods, another important future area of research will involve detailed and large scale benchmarking for various problems.
We also plan to demonstrate that the use of CTQWs on problem specific graphs provides significant advantage for non-binary problems, such as the quadratic assignment problem, over the alternative approach of converting these to QUBO problems and making use of the standard QAOA/binary mixer.